Find the sum Σ, (3: +1)
Answer #1

Σ(3x + 1) k=1 = (3 - 1 + 1) + (3 - 2+1) + (3 - 3 + 1) + (3 - 4+1) = (3 + 1) + (6 + 1) + (9 + 1) + (12 +1) = 4+7+ 10 + 13 = 34

Alternative Method: (3k + 1) k= 1 k1 n(n + 1) = 3.4+ 1) +1x1 [ * = n(n+1) = 3. 4(4+1) 2+1x4 2 Lk=1 = 32.5 +4 = 30 +4. = 34